Complete combustion of 8.90 g of a hydrocarbon produced 28.4 g of CO2 and 10.2 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.


1
Expert's answer
2020-10-14T14:01:00-0400

First, find the mass of carbon in an unknown hydrocarbon:

m(C) = 12*28.4/44 = 7.745 g

Now we find the mass of hydrogen in the unknown hydrocarbon:

m(H) = 2*10.2/18 ​= 1.133 g

Now we find the ratio of the amount of carbon and hydrogen in an unknown substance:

n(C):n(H) = (7.745/12):(1.133/1) = 0.645:1.133 = 1:1.75

We found the ratio of carbon and hydrogen atoms, but the values are not true. Therefore, we multiply each number by 8, so that we get integer values:

n(C):n(H) = 8:14

The formula of compound is C8H14.